Find the odds in favor of rolling the same number on a pair of dice.: Find the odds in favor of rolling the same number on a pair of dice.
Answer by jim_thompson5910(9376)

There are 36 different combinations that you can roll with a pair of dice (since 6*6=36). If you aren't convinced, you can make a table to diagram all of the possible outcomes.

Now let's list the pairs of numbers that are identical:

1,1
2,2
3,3
4,4
5,5
6,6


Since there are 6 choices, this means that the probability of rolling the same number on a pair of dice is 6/36=1/6